# VitaGlyph: Vitalizing Artistic Typography with Flexible Dual-branch Diffusion Models

### v2

@@ H Additional Results
We provide more Chinese artistic typography examples in Fig. J and examples with multiple customized concepts in Fig, L. In Fig. L, we can observe that we incorporate “coffee beans” and “doughnut” in artistic typography of “Starbucks”, with its texture style is “green and white background”. And we can also generate “Hello World” in Chinese, Japanese, Korean and English with concepts like “Robot” and “The Earth”.
I Limitations
@@
Fig. L’s “Jordan” shows that our method would exhibit some deficiencies when generating humans. We believe this is due to the base model, StableDiffusion-1.5. In future work, we will address and refine these issues. Additionally, while we utilize multiple parallel diffusion models for compositional generation, incorporating more than two concepts with different models poses challenges in terms of memory usage and computation time and we will optimize our implementation as soon as possible.
J Future Work
@@
In addition to addressing the limitations mentioned above, there exist several promising avenues for future research. [26, 46] delve into the animation of artistic words. Offering insights into the harmonious integration of movement and typography artistry, these studies pave the way for us to explore into the dynamic and kinetic dimensions of typographic expression. Additionally, Our approach are the first to engage artistic typography with compositional generation. We are honored to develop a artistic typography dataset and improve our benchmark with more methods and metrics. This provides a foundation for the development of innovative methods to enhance artistic typography techniques.
K Social Impacts
@@
VitaGlyph has the potential to significantly impact the creative industry by offering new avenues for designers and artists to explore innovative and appealing typographic designs. This technology not only enhances artistic expression and readability but also has the capacity to elevate user experiences in fields such as advertising, branding, and digital media. Moreover, VitaGlyph’s interdisciplinary approach, combining text generation, image processing, and typography design, can foster cross-domain collaborations, driving technological advancements and fostering innovation. Without proper guidance and ethical considerations, individuals may engage in harmful practices such as image forgery or digital impersonation. We advocate for the development of legal frameworks that address AI-generated content, including penalties for malicious use.
Figure L: More examples of VitaGlyph’s application. Our method can be applied to various languages and fonts. With multiple user-specific concepts injected and the utilization of Controllable Compositional Generation, we manage to design logo or slogan like “CVPR”, “VitaGlyph” and “Hello World”..

### v3

@@ IX Additional Results
We provide more examples with multiple customized concepts in Fig. A. We can observe that we incorporate “the Earth”, “doughnut” and “robot waving hands” in artistic typography of “Hello World”, with its texture style is “colorful candy texture”. And we can also generate “Hello World” in Chinese, Japanese, Korean, and English.
X Limitations
@@
As shown in figure 6, our method would exhibit some deficiencies when generating humans. We believe this is due to the base model, StableDiffusion-1.5. In future work, we will address and refine these issues. Additionally, while we utilize multiple parallel diffusion models for compositional generation, incorporating more than two concepts with different models poses challenges in terms of memory usage and computation time, and we will optimize our implementation as soon as possible.
XI Future Work
@@
In addition to addressing the limitations mentioned above, there exist several promising avenues for future research. (Liu et al. 2024b; Xie et al. 2023) delve into the animation of artistic words. Offering insights into the harmonious integration of movement and typography artistry, these studies pave the way for us to explore the dynamic and kinetic dimensions of typographic expression. Additionally, Our approach is the first to engage artistic typography with compositional generation. We are honored to develop a artistic typography dataset and improve our benchmark with more methods and metrics. This provides a foundation for the development of innovative methods to enhance artistic typography techniques.
XII Social Impacts
@@
VitaGlyph has the potential to significantly impact the creative industry by offering new avenues for designers and artists to explore innovative and appealing typographic designs. This technology not only enhances artistic expression and readability, but also has the capacity to enhance user experiences in fields such as advertising, branding, and digital media. Moreover, VitaGlyph’s interdisciplinary approach, combining text generation, image processing, and typography design, can foster cross-domain collaborations, driving technological advancements, and fostering innovation. Without proper guidance and ethical considerations, people can engage in harmful practices such as image forgery or digital impersonation. We advocate for the development of legal frameworks that address AI-generated content, including penalties for malicious use.
XIII Time Efficiency
Model Time(Sec.)
Word-as-image (Iluz et al. 2023) 102
WordArt Designer (He et al. 2023) 116
VitaGlyph (ours) 15
Table C: We computed the average time required to generate each image over the course of generating 10 images of the same size on a NVIDIA Tesla V100.
Since VitaGlyph operates entirely within the raster domain for both transformation and rendering, it differs from WordArt Designer and Word-as-image, which utilize vector-based methods with SDS Loss (Poole et al. 2022) for glyph deformation. This distinction allows us to significantly reduce the inference time, as demonstrated in Tab. C.